OpenStudy (anonymous):

Find the solutions in the interval [0,2pi) 2cos3theta=1

OpenStudy (anonymous):

is the problem \[2\cos ^{3}\theta=1\] or \[2\cos(3\theta)=1\]

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

ok thanks..

OpenStudy (anonymous):

cos(3x)=1/2 \[3x=2n \pi \pm \pi/9\]

OpenStudy (anonymous):

ok, use what not so bright said but remember to keep within [0, 2pi)

OpenStudy (anonymous):

I don't understand what he put.. but I got theta = pi/6 , 2pi/6 is that right?

OpenStudy (anonymous):

no @sportscrazy12

OpenStudy (anonymous):

okay.. well I don't know how do this problem

OpenStudy (anonymous):

answers are obtained by putting n=1 ,2,3

OpenStudy (anonymous):

but where did you get pi/9?

OpenStudy (anonymous):

put n=0 as well

OpenStudy (anonymous):

well its only looking for the the interval between [0, 2pi) I don't think I'd need to add integers to 2pi

OpenStudy (anonymous):

r u there?

OpenStudy (anonymous):

yupp

OpenStudy (anonymous):

 forget about the problem a bit. what angle should be in place of the ?